INFRARED SPECTRA OF CO_{2} - H_{2} COMPLEXES

\underline{A.R.W. McKELLAR}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

Infrared spectra of weakly-bound CO_{2}-H_{2} complexes have been studied in the region of the CO_{2} ν_{3} asymmetric stretch, using a tunable diode laser probe and a pulsed supersonic jet expansion. For CO_{2}-paraH_{2}, results were obtained for three isotopic species, ^{12}C^{16}O_{2}, ^{13}C^{16}O_{2}, and ^{12}C^{18}O_{2}. These spectra were analyzed using an asymmetric rotor Hamiltonian, with results that resembled those obtained previously for OCS- and N_{2}O-paraH_{2} [1], except that half the rotational levels were missing due to the symmetry of CO_{2} and the spin statistics of the ^{16}O or ^{18}O nuclei (as in the similar case of CO_{2}-He [2]). The CO_{2}-paraH_{2} complex has a T-shaped structure with an intermolecular distance of about 3.5 Å, and the CO_{2} ν_{3} vibration exhibits a small red shift (-0.20 cm^{−1}) in the complex. For CO_{2}-orthoH_{2}, more complicated spectra were observed which could not be assigned, in contrast to OCS- and N_{2}O-H_{2} where the paraH_{2} and orthoH_{2} spectra were similar, though distinct.

[1] J. Tang and A.R.W. McKellar, J. Chem. Phys. 116, 646 (2002); 117, 8308 (2002).

[2] M.J. Weida, J.M. Sperhac, D.J. Nesbitt, and J.M. Hutson, J. Chem. Phys. 101, 8351 (1994).